Verify that the given functions form the fundamental set of solution of the differential equation on the given indicated interval.

y" — 7y' + 10y = 24e^x
y = Cl1e^2x + C2e^5x + 6e^x, (-0, infinity)
1
Expert's answer
2021-09-28T01:21:23-0400

Solution;

The symbolic form of the equation is;

"(D^2-7D+10)y=24e^x"

The auxiliary equation is;

"m^2-7m+10=0"

Rewrite as the following;

"m^2-2m-5m+10=0"

"m(m-2)-5(m-2)=0"

"(m-2)(m-5)=0"

m=2 or 5

Hence;

"C.F=y_c" ="C_1e^{2x}+C_2e^{5x}"

To find the particular solution;

Take,

"y_1=e^{2x}" ,"y_2=e^{5x}" and "X=24e^x"

Find the Workisian;

"W=\begin{bmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{bmatrix}" ="\begin{bmatrix}
 e^{2x} & e^{5x} \\
 2e^{2x} & 5e^{5x}
\end{bmatrix}"

"W=5e^{7x}-2e^{7x}=3e^{7x}"

"P.I=y_p=uy_1+vy_2"

Take ;

"u=-\int\frac{y_2X}{W}=-\int\frac{e^{5x}24e^x}{3e^{7x}}=8e^{-x}"

"v=\int\frac{y_1X}{W}=\int\frac{e^{2x}24e^x}{3e^{7x}}=-2e^{-4x}"

"P.I=y_p=8e^{-x}e^{2x}+-2e^{-4x}e^{5x}"

"y_p=8e^x-2e^x=6e^x"

The general solution is;

"y=C.F+P.I=y_c+y_p"

"y=C_1e^{2x}+C_2e^{5x}+6e^x"

Hence the solution is verified.